Find the slope of a line perpendicular to a line passing through (5,6) and (-9,10)
Answer by saifoo.khan(162)

Slope of this line = (y2-y1)/(x2-x1)
(5,6) and (-9,10)
(10-6)/(-9-5)
4/-14
-2/7
The slope of this current line is -2/7
now to find the slope of the perp line, this formula is used
m1 * m2 = -1
slope of first line times slope of perpendicular = -1
so,
-2/7 * m2 = -1
2/7 * m2 = 1
m2 = 7/2
So now u have the slope of it's perpendicular line,
Hope u got the answer required.
====================================
If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($15/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($1.5-$3.0 per problem). Contact me at einstein.saifoo@gmail.com
Thank you. =)